12/(x^2-4)-3/(x-2)=5/(x+2)
Answer by stanbon(57219)

12/(x^2-4)-3/(x-2)=5/(x+2)
-----
Multiply thru by the lcm: lcm: x^2-4
-----
12 - 3(x+2) = 5(x-2)
12 - 3x - 6 = 5x - 10
8x = 16
x = 2
But if x = 2, x^2-4 = 0.
So, there are no solutions for the original problem.
=============
Cheers,
Stan H.